OpenStudy (anonymous):

Find the first triangular number which is greater than 1000????

OpenStudy (amistre64):

1,3,7,13 ... those things right?

OpenStudy (anonymous):

1,3,6,10

OpenStudy (amistre64):

yeah, 1,3,6,10 :)

OpenStudy (amistre64):

pascals triangle might be useful

OpenStudy (anonymous):

I know it is be some sort of inequality.

OpenStudy (anonymous):

I don't know what pascals triangle is.

OpenStudy (anonymous):

Its all the way at the back of my text.

OpenStudy (amistre64):

1 1 1 1 2 1 1 3 3 1 1 4 6 4 1 1 5 10 10 5 1 the 3rd inset is our trinagle numbers nC2

OpenStudy (amistre64):

n(n-1) ------ > 1000 2 n^2 -n > 2000

OpenStudy (amistre64):

n^2 -n -2000 > 0

OpenStudy (amistre64):

that equals 0 when n=45.xxx; so 46 is my bet

OpenStudy (anonymous):

45 my text says.

OpenStudy (amistre64):

45C2 = 45(22) = 990

OpenStudy (amistre64):

the triangles start with 2C2 so yes; 45 is it :/

OpenStudy (amistre64):

another method i recall has to do with finding the degree of differences 1 3 6 10 15 2 3 4 5 1 1 1 and the first term in each row is the coeff of a factorial exapnasion for lack of a more descriptive term 1 + 2n + 1n(n-1) 1 + 2n + n^2 -n n^2 +n + 1 but im sure im remembering it a little off tho

OpenStudy (anonymous):

>>> i = 1 >>> n = i >>> while n < 1000: ... i += 1 ... n += i ... >>> i 45

OpenStudy (anonymous):

Python

OpenStudy (amistre64):

\[t_n=t_{n-1}+n;\ t_0=0\] \[t_n=t_{n-2}+(n-1)+n\] \[t_n=t_{n-3}+(n-2)+(n-1)+n\] \[t_n=t_{n-4}+(n-3)+(n-2)+(n-1)+n\] \[t_n=t_{n-r}+(n-(r-1))+(n-(r-2))+(n-(r-3))+...+(n-3)+(n-2)+(n-1)+n\] when r=n \[t_n=t_{0}+(1+2+3+...+(n-3)+(n-2)+(n-1)+n\] \[t_n=\frac{n(n+1)}{2}\] \[\frac{n(n+1)}{2}>1000\] \[n(n+1)>2000\] \[n^2+n-2000>0\] \[n>\frac{-1+\sqrt{8001}}{2}::\frac{88.5}{2}\to\ 44.25\]

OpenStudy (amistre64):

.... the antiPython lol

OpenStudy (anonymous):

Yes, in some ways in makes me feel dirty. But it's fast and easy.